Suppose the market of carpets is competitive. The demand for and the supply of carpets in the market have been estimated as follows: Demand: Qd = 6500 – 100P Supply: Qs = 1200P
A typical firm producing carpets has a total cost function of C = 100 +
1
Expert's answer
2019-12-09T09:54:47-0500

 Given,"Qd= 6500- 100P," "Qs= 1200P"

 At equilibrium "Qd=Qs"

Therefore, 6500-100P= 1200P.

1300P= 6500

P= 6500/1300= 5

Qs= 1200P = 1200*5 = 6000

Qd =6500-100*5 = 6000